OpenStudy (anonymous):

Can you show me how to work out this problem? Apparently, the answer is 7.32%, but I can't seem to simplify it. .06 × (1 - .30) ($1,500,000 + .12 × $1,000,000) / ($1,500,000 + $1,000,000)

OpenStudy (anonymous):

can you show the steps you've gone thru? You need to pay attention to the order of operations - so for example you do the .12 x 1 million BEFORE you add the 1.5 million to it. so it SHOULD be written like this: .06 × (1 - .30) ($1,500,000 + (.12 × $1,000,000)) / ($1,500,000 + $1,000,000) so that that order of operations is obvious. Show your steps and so we can find a problem.

OpenStudy (anonymous):

Parenthesis, Exponents, Multiplication, Division, Addition, Subtraction, so... Parenthesis: .12x1mill = 120,000 Then, 120,000 + 1,500,000 = 1,620,000 So far, so good?

OpenStudy (anonymous):

Then, 1-.30 = .70 Then, .70 x 1,620,000 = 1,134,000 Then, 1,134,000 x .06 = 68040 ( I keep getting this number) But when I divide the bottom number ($1,500,000 + $1,000,000) = 2,500,000 = 68040 / 2,500,000 doesn't = 7.32% or anything close to it. I get, .027216

OpenStudy (anonymous):

salam22- sorry for the delay- i get the same result you do- so i can only suggest that the question is wrong, and you need to bring it up with your instructor. Does it have an actual problem associated with it, or was it just the math? I'm guessing that it's something to do with the precedence.